OpenStudy (anonymous):

sq root of x + 10 = x -2

OpenStudy (anonymous):

\[\sqrt{x+10}=x-2\] Is this the equation? Are you looking for the value of x?

OpenStudy (anonymous):

yes

hartnn (hartnn):

the first step is to square both sides,

OpenStudy (anonymous):

\[\sqrt{x+10}=x-2\] \[(\sqrt{x+10})^2=(x-2)^2\] \[x+10=x^2+4\] \[x+6=x^2\] You should be able to solve the rest now

hartnn (hartnn):

\[(x-2)^2=x^2-4x+4\]

hartnn (hartnn):

\[x+10 = x^2-4x+4\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Yes sorry misread that hartnn has the last step

OpenStudy (anonymous):

so then you get \[x-6=x^2-4x\] and solve from there

hartnn (hartnn):

\[x+6=x^2-4x\]

hartnn (hartnn):

\[x^2-5x-6=0\] can you solve this quadratic ?